Using identities, evaluate. 
5.2^2 ???????????????????????

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

5.2^2= (5 + 0.2)^2
= 5^2+ 2 + 0.2^2
= 25 + 2 + 0.4
=27.4

Thanks